A wire or length L is made into a semicircle or a half circle. Find Moment of inertia passing through any of its ends

Grade:11

1 Answers

Vikas TU
7 years ago
About one of it’s end,
I = Icom + Md^2 where d = 2R/pi is the radius of ring.
Now for Icom = I – Md^2
Icom = Mr^2(1 – 4/pi^2)
 
and from pythodorus theorem,
h^2 let = R^2[1 + 4/pi^2]
 
Hence,
Total monet of inertia =>
 
 R^2[1 + 4/pi^2] +  R^2[1 - 4/pi^2]